Barnes Company reports the following operating results for the month of August: sales $325,000 (units 5,000); variable costs $213,000; and fixed costs $71,600. Management is considering the following independent courses of action to increase net income.

Compute the net income to be earned under each alternative.

1. Increase selling price by 10% with no change in total variable costs or sales volume.

Net income $

Entry field with incorrect answer now contains modified data


2. Reduce variable costs to 58% of sales.

Net income $

Entry field with incorrect answer now contains modified data


3. Reduce fixed costs by $24,000.

Net income $

Entry field with incorrect answer now contains modified data
Answer #1

2 Selling price 3 Increases selling price 4 Sales 5 Variable costs 6 Fixed costs 7 Net income - 1 1 2 3 $ 65 $ 71.50 $ 357,50

2 Selling price =325000/5000 3 Increases selling price =+B2*110% 4 Sales =5000*B3 5 Variable costs 213000 6 Fixed costs 71600